OpenStudy (xxcookiequeenxx):

What is the missing value? 3^6 ----=3^4 3^--- A. −10 B. −2 C. 2 D. 10

OpenStudy (xxcookiequeenxx):

@AllNamesHaveBeenTaken

hartnn (hartnn):

\(\Large \dfrac{x^n}{x^m}= x^{n-m}\)

OpenStudy (faiqraees):

\[\large\rm \frac{a^x}{a^y} = a^{x-y} \]

hartnn (hartnn):

\(\Large \dfrac{3^6}{3^x} = 3^{6-x} = 3^4\) so comparing the exponents, 6-x =4 x =... ?

OpenStudy (xxcookiequeenxx):



hartnn (hartnn):

is it -6 or +6 ??

OpenStudy (xxcookiequeenxx):

-6

hartnn (hartnn):

ok, then -6-x = 4 get x